OpenStudy (anonymous):

two standard dice are rolled. Let A be the event that the first die lands on a number larger than 4. Let B be the event that the second die lands on a number larger than 2. Which of the following are true. Check all that apply. A. P(B)= 1/6 B. P(A and B)= 1/36 C. P(A and B)= 1/6 D. P(A)= 1/2

OpenStudy (campbell_st):

P(A) = 2/6 or 1/3 P(B) = 3/6 or 1/2 P(A and B) = P(A) x P(B) hope this helps

OpenStudy (anonymous):

Okay thank you!